OpenStudy (anonymous):

-9 divided by (- 4) (show as an improper fraction)

ganeshie8 (ganeshie8):

how many (-4)'s are there in a (-9) ?

ganeshie8 (ganeshie8):

\[\large \begin{align} \dfrac{(-9)}{(-4)} &= \dfrac{(-4) + (-4)+(-1)}{(-4)}\\~\\ &= \dfrac{(-4)}{(-4)}+\dfrac{(-4)}{(-4)}+\dfrac{(-1)}{(-4)}\\~\\ &= 1+1+\dfrac{(-1)}{(-4)}\\~\\ &=2+\dfrac{(-1)}{(-4)}\\~\\ &=2\dfrac{(-1)}{(-4)}\\~\\ \end{align}\]

OpenStudy (coconutjj):

9/4

OpenStudy (coconutjj):